Bond Discount, Entries for Bonds Payable Transactions, Interest Method of Amortizing Bond Discount

On July 1, 20Y1, Livingston Corporation, a wholesaler of manufacturing equipment, issued $63,000,000 of 20-year, 11% bonds at a market (effective) interest rate of 14%, receiving cash of $50,403,780. Interest on the bonds is payable semiannually on December 31 and June 30. The fiscal year of the company is the calendar year.

Required:

For all journal entries, if an amount box does not require an entry, leave it blank.

1. Journalize the entry to record the amount of cash proceeds from the issuance of the bonds.

20Y1 July 1

2. Journalize the entries to record the following:

a. The first semiannual interest payment on December 31, 20Y1, and the amortization of the bond discount, using the interest method. Round to the nearest dollar.

20Y1 Dec. 31

b. The interest payment on June 30, 20Y2, and the amortization of the bond discount, using the interest method. Round to the nearest dollar.

20Y2 June 30

3. Determine the total interest expense for 20Y1. Round to the nearest dollar.
$
Answer #1
1
20Y1 July 1 Cash 50403780
Discount on Bonds payable 12596220
       Bonds payable 63000000
2
a
20Y1 Dec. 31 Interest expense 3528265 =50403780*14%*6/12
       Discount on Bonds payable 63265
       Cash 3465000 =63000000*11%*6/12
b
20Y2 June 30 Interest expense 3532693 =(50403780+63265)*14%*6/12
       Discount on Bonds payable 67693
       Cash 3465000
3
Total interest expense for 20Y1 3528265